An agricultural economist estimated the following linear demand function for milk in Queensland: Q = 857 - 18.1P where Q represents the number of two litre milk bottles demanded per day in thousands and P represents the price of these bottles in dollars. If on a given day the two litre milk bottle retails for $3.25, predict the number of two litre milk bottles demanded in thousands (to nearest whole thousand).
Answer #1

in , Q = 857-18-19 =857 - 18.1($3.25 - 798.175 ? 798
Kindly upvote:)